# Integrated Multimodal Sensing and Communication: Challenges, Technologies, and Architectures

### v1

@@ Abstract
The evolution towards 6G networks requires the intelligent integration of communication and sensing capabilities to support diverse and complex applications, such as autonomous driving and immersive services. However, existing integrated sensing and communication (ISAC) systems predominantly rely on single-modal sensors as primary participants, which leads to a limited representation of environmental features and significant performance bottlenecks under the emerging requirements of 6G applications. This limitation motivates a paradigm shift from single-modal to multimodal ISAC. In this article, we first analyze the key challenges in realizing multimodal ISAC, including the fusion of heterogeneous multimodal data, the high communication overhead among distributed sensors, and the design of efficient and scalable system architectures. We then introduce several enabling technologies, such as large AI models, semantic communication, and multi-agent systems, that hold promise for addressing these challenges. To operationalize these technologies, we zoom into three architectural paradigms: fusion-based multimodal ISAC (F-MAC), interaction-based multimodal ISAC (I-MAC), and relay-based multimodal ISAC (R-MAC), each tailored to organize devices and modalities for efficient collaboration in different scenarios. Thereafter, a case study is presented based on the F-MAC scheme, demonstrating that the scheme achieves more comprehensive sensing and improves sensing accuracy by approximately 80% compared to conventional single-modal ISAC systems. Finally, we discuss several open issues to be addressed in the future.
@@ I-A Background
Integrated sensing and communication (ISAC), as an emerging paradigm, unifies the functions of wireless sensing and communication within a shared hardware and spectral framework [1]. By leveraging common radio frequency (RF) signals for both tasks, ISAC enables efficient resource utilization, reduces hardware redundancy, and enhances spectral efficiency, making it a promising solution for future intelligent wireless systems. However, traditional ISAC approaches predominantly rely on RF signals as the sole modality for sensing, which limits the system’s capability in complex environments [2]. RF signals provide a single-modality perspective, offering limited information, such as range, velocity, and coarse spatial features, which is insufficient for capturing fine-grained semantic or contextual details of complex environments. Additionally, RF-based sensing often struggles to adapt to rapidly changing or highly dynamic environments, as its performance may degrade significantly due to variations in multipath propagation, interference, and occlusions. These limitations promote the paradigm shift from single-modal to multimodal ISAC.
Different types of sensors operate on fundamentally distinct principles and collect sensory data with varying properties. Examples of these sensing techniques include radar, light detection and ranging (LiDAR), red-green-blue-depth (RGB-D) cameras, and the global positioning system (GPS), which are collectively referred to as multimodal sensing. By leveraging the complementary strengths of different modalities, multimodal ISAC holds great promise in enhancing sensing accuracy, robustness, and generalization. For instance, the inclusion of multimodal sensing can accelerate the beam selection process, such as LiDAR-aided beam prediction [3], camera-GPS-aided beam prediction [4], and camera-aided beam prediction [5], where conventional methods require an exhaustive search across all candidate beam pairs. Furthermore, multimodal ISAC systems demonstrate improved adaptability to dynamic scenarios. For example, visual data can compensate for RF degradation under adverse conditions, such as multipath fading or obstruction.
@@ I-B Contributions
    Architectures for multimodal ISAC: To systematically integrate these technologies, we examine three distinct architectural paradigms tailored to different multimodal collaboration patterns: (i) Fusion-based Multimodal ISAC (F-MAC), which emphasizes centralized semantic fusion; (ii) Interaction-based Multimodal ISAC (I-MAC), which enables direct peer-to-peer semantic interaction among sensors; and (iii) Relay-based Multimodal ISAC (R-MAC), which focuses on the large range and long distance of sensing tasks across modalities. These frameworks provide flexible and scalable design options for deploying ISAC in diverse application environments.
@@
    Case Validation: We conduct a case study based on the F-MAC architecture to validate the framework. Experimental results demonstrate that the F-MAC scheme achieves enhanced perceptual coverage and improves overall sensing accuracy by approximately 80% compared to conventional single-modal ISAC approaches, thereby showcasing the practical benefits of multimodal integration.
@@ II-A Heterogeneous Multimodal Fusion
Multimodal ISAC inherently involves the integration of data streams originating from diverse sensing modalities such as RF signals, vision sensors, and LiDAR. These modalities differ substantially in terms of data dimensionality, spatial-temporal resolution, coverage, and semantic abstraction. For instance, aligning 2D visual images with 1D radar waveforms is particularly challenging due to discrepancies in data structure, coordinate systems, and information content. Moreover, inconsistencies in sensing frequency and perception latency further complicate synchronized fusion.
@@ II-B High Communication Overhead
Real-time sensing and decision-making in distributed ISAC systems require frequent and high-volume data exchanges between sensors, edge nodes, and centralized servers. This results in considerable bandwidth consumption and elevated energy expenditure, particularly in wireless or resource-constrained environments. The issue becomes more pronounced in low-dynamic scenarios—such as static nighttime surveillance—where redundant or minimally informative data (e.g., successive identical video frames) continue to be transmitted, thereby wasting transmission resources without improving situational awareness.
@@ II-C Context-Aware Architectural Design
Designing a system architecture that can adapt to diverse operational environments and task requirements remains a fundamental challenge. Rigid deployment modes may offer simplicity in control and resource management but often lack the flexibility needed to accommodate the heterogeneous demands of ISAC applications. For example, while centralized architectures benefit from powerful computing infrastructure, they introduce latency and dependency issues in time-critical or infrastructure-sparse settings. Conversely, purely distributed systems offer autonomy but often suffer from limited coordination and reduced global situational awareness.
@@ III-C Multi-Agent System
MAS [8] provides the structural foundation for enabling scalable, decentralized intelligence within the multimodal ISAC. In MAS, multiple autonomous agents, either software-based or embodied in hardware, interact with the environment, as well as each other, to collaboratively achieve task objectives. Each agent perceives local conditions, makes decisions, and executes actions, often through cooperative reasoning and intent-aware communication. As shown in Fig. 1(c), the core component of MAS includes four functional modules: a knowledge base for storing domain knowledge, a memory module for accumulating interaction history, a large language model (LLM), such as Deepseek and GPT, for reasoning and decision-making, and external tools for executing actions. During operation, an agent continuously interacts with the environment, stores relevant observations in memory, combines them with contextual knowledge from the knowledge base, and inputs the fused information into the LLM. The LLM then analyzes the situation and invokes appropriate tools (e.g., an AI or generic algorithm) to perform task-specific actions.
@@
Figure 2: The illustration of the proposed three architectural schemes of the multimodal ISAC. (a) Fusion-based Multimodal ISAC. (b) Interaction-based Multimodal ISAC. (c) Relay-based multimodal ISAC.
IV Architectures for Multi-Modal ISAC
As illustrated in Fig. 2, this section presents three architectural schemes tailored to distinct scenario-specific requirements, in which the identified technologies, LAM, SC, and MAS, are selectively and strategically integrated to enhance both the efficiency and adaptability of the multimodal ISAC.
IV-A Fusion-based Multimodal ISAC
@@
As illustrated in Fig. 2(a), the F-MAC scheme adopts a centralized fusion architecture to integrate semantic information from heterogeneous sensors. Specifically, each sensor is paired with a lightweight edge agent. These edge agents function as semantic encoders, responsible for extracting and compressing task-relevant semantics from local multimodal data. Then, the resulting compact semantic representations are transmitted to a base station for further processing over a wireless channel. At the base station, LAMs serve as a central agent to align and fuse the received multimodal semantics to generate unified representations. The fused features can then be utilized for downstream applications such as global decision-making, collaborative sensing, and environmental modeling. This process is detailed in Fig. 1(a) and Fig. 1(b). A typical application scenario of this solution is intelligent transportation, where cameras, LiDAR sensors, and RF devices are deployed across various intersections [9]. Each sensor, via its corresponding edge agent, extracts semantic information such as vehicle positions, motion trajectories, and Doppler shifts. The compressed semantic features are transmitted to a central LAM located at a traffic control center, where global fusion is performed to achieve an accurate understanding of real-time traffic conditions, thereby enabling intelligent traffic signal coordination and congestion mitigation.
@@
In the F-MAC scheme, the use of a powerful LAM at the central agent enables deep reasoning over globally aggregated data, improving decision accuracy. Meanwhile, SC offloads bandwidth-intensive raw data transmission from the edge and ensures efficiency in bandwidth-limited environments. However, this scheme is heavily reliant on the computational capabilities of the central agent. Consequently, when communication links between edge sensors and the central base station are disrupted, edge agents are unable to operate autonomously. This dependency can lead to reduced system responsiveness and efficiency, particularly in dynamic scenarios with intermittent connectivity.
IV-B Interaction-based Multimodal ISAC
Fig. 2(b) illustrates the I-MAC scheme, which aims to decentralize intelligence across the network fully. In this scheme, each sensor is equipped with a language model-driven edge agent, as depicted in Fig. 1(c). Each edge agent independently processes its multimodal sensory inputs and transmits only essential high-level semantic outcomes, such as alerts, decisions, or event summaries, to neighboring agents. Subsequently, each agent performs local analysis, reasoning, and task-specific decision-making by leveraging the received semantic information, internal memory, and localized knowledge bases. A representative application scenario is industrial inspection, where cameras, vibration sensors, and acoustic sensors are deployed along a production line [10]. Each sensor, empowered by an AI agent, semantically interprets its sensory stream. For instance, if one agent detects an abnormal vibration and another simultaneously identifies a visual defect, they can directly exchange semantic messages to collaboratively diagnose and localize faults in real time, without involving a remote processing center.
The I-MAC architecture provides maximum autonomy, scalability, and resilience. By minimizing reliance on centralized network infrastructure, it is particularly effective in bandwidth-constrained, delay-sensitive, or infrastructure-deficient environments. More importantly, compared to the F-MAC scheme, edge agents in I-MAC are not limited to passive data providing; they act as independent, task-aware agents with cognitive and communicative capabilities. Nonetheless, data heterogeneity and differences in prior knowledge across edge agents may lead to cognitive inconsistencies. For example, under low-light conditions, a camera-based agent may struggle to accurately perceive environmental information, whereas an infrared-based agent can operate normally. In such cases, the agents may generate divergent semantic interpretations of the same scene, potentially leading to conflicting local decisions and degraded overall system performance. Furthermore, due to hardware limitations, resource-constrained edge sensors may only support lightweight or compressed language models, which can introduce trade-offs in inference accuracy and reasoning capability.
IV-C Relay-based multimodal ISAC
Fig. 2(c) illustrates the R-MAC scheme, where each sensor is equipped with a moderately capable edge agent responsible for local feature extraction and task-relevant semantic interpretation. These edge agents extract essential information from raw multimodal data and transmit it to nearby relay nodes, such as road side units (RSU), for further processing. The relay agents then interpret, repackage, and forward the semantic information to other edge agents operating on different modality sensors. Finally, the receiving edge agent performs semantic analysis and makes task-specific decisions based on the parsed information. For example, in the drone surveillance scenario, a camera-equipped drone employs its onboard edge agent to extract positional semantics from aerial observations. This semantic information is transmitted to an RSU-based relay agent, which references a lightweight knowledge base. If the observed object is unrecognized, the relay agent formulates a semantic directive (e.g., “focus radar on coordinates (x, y)”) and forwards it to a radar sensor positioned on the ground. Upon receiving the directive, the radar’s edge agent dynamically adjusts its ISAC beam toward the specified coordinates, thus avoiding exhaustive scanning and improving sensing efficiency.
Compared to the F-MAC scheme, R-MAC supports longer distances for sensing tasks. Compared to F-MAC, it supports extended spatial coverage and reduces dependency on a central fusion node. However, the effectiveness of R-MAC depends heavily on robust SC protocols and accurate message translation between modalities. Additionally, it may underperform in scenarios that require either fine-grained, globally consistent fusion (as enabled by F-MAC) or high levels of autonomous decision-making at the edge (as in I-MAC).
To clearly illustrate the differences among the three schemes, Table I provides a comparative summary of their key characteristics. Overall, F-MAC is ideal for infrastructure-rich environments requiring centralized fusion and global decision-making (e.g., intelligent transportation hubs and smart cities). I-MAC suits infrastructure-sparse or bandwidth-limited scenarios where autonomous, peer-to-peer interaction is essential (e.g., industrial inspection sites). R-MAC bridges the two, supporting cross-modal coordination in semi-structured settings with moderate infrastructure and constrained edge devices (e.g., drone surveillance and edge patrolling systems).
@@ IV-C Relay-based multimodal ISAC
Aspect F-MAC I-MAC R-MAC
Strengths High-accuracy inference; Strong multimodal fusion High autonomy; Strong scalability; Robust resilience Balanced workload; Moderate communication efficiency
Weaknesses Strong central dependency; Low fault tolerance Cognitive inconsistency; High edge computational burden Protocol dependency; Compromised global/local reasoning
Communication Load High Low Moderate
Fault Tolerance Low (single point of failure) High (fully decentralized) Moderate (relay-dependent)
Key Technologies LAM and SC MAS and SC SC
Use Cases Intelligent transportation hubs Industrial inspection sites Drone surveillance

### v2

# A Unified Cloud-Edge-Terminal Framework for Multimodal Integrated Sensing and Communication

@@ Abstract
The transition to 6G calls for tightly integrated sensing and communication to support mission-critical services such as autonomous driving, embodied AI, and high-precision telemedicine. However, most existing ISAC designs rely on a single sensing modality (often RF), which limits environmental understanding and becomes a bottleneck in complex and dynamic scenes. This motivates a shift from single-modal to multimodal ISAC, where heterogeneous sensors (e.g., radar, LiDAR, and cameras) complement each other to improve robustness and semantic awareness. In this article, we first summarize key challenges for multimodal ISAC, including heterogeneous fusion, communication overhead, and scalable system design. We then highlight three enabling technologies: large AI models, semantic communications, and multi-agent systems, and discuss how their combination can enable task-oriented multimodal perception. Building on these insights, we propose a unified cloud–edge–terminal (CET) framework that hierarchically distributes intelligence and supports three adaptive operation modes: global fusion mode (GFM), cooperative relay mode (CRM), and peer interaction mode (PIM). A case study evaluates the framework across three modes, demonstrating that GFM achieves the highest accuracy, PIM minimizes latency, and CRM strikes an optimal balance between performance and efficiency. Finally, we conclude with open research issues and future directions.
@@ I-A Background
Next-generation intelligent applications, such as autonomous driving, embodied artificial intelligence (AI), and high-precision telemedicine, inherently rely on the comprehensive environmental understanding and ultra-reliable connectivity [4]. For instance, embodied robots require visual semantic understanding for navigation, whereas remote surgical manipulators demand haptic and visual feedback coupled with low-latency control. This underscores the critical role of integrated sensing and communication (ISAC) in future network architectures.
Although traditional ISAC systems effectively unify sensing and communication within a shared radio frequency (RF) framework to enhance spectral efficiency, relying exclusively on RF signals as the sole sensing modality often results in significant performance bottlenecks [13]. While RF signals are proficient in range and velocity estimation, they offer a restricted single-modality perspective that is often insufficient for capturing the fine-grained semantic and contextual details demanded by complex environments. Moreover, RF-based sensing is susceptible to severe performance degradation in highly dynamic scenarios caused by multipath propagation, interference, and occlusions.
These limitations necessitate a fundamental paradigm shift from single-modal to multimodal ISAC. In contrast to single-modal approaches, multimodal ISAC integrates data from diverse sensors operating on distinct physical principles, such as radar, Light Detection and Ranging (LiDAR), RGB-D cameras, and Global Positioning System (GPS), to construct a holistic representation of the physical world. By leveraging the complementary strengths of these diverse modalities, multimodal ISAC offers significant potential for enhancing sensing accuracy, robustness, and generalization capabilities [6]. Existing research has already demonstrated the efficacy of this synergy; for instance, incorporating visual or spatial modalities can significantly accelerate the beam selection process. Notable implementations include LiDAR-aided [9], camera-GPS-aided [3], and camera-aided beam prediction [2], where multimodal data mitigates the latency associated with the exhaustive search protocols typical of conventional RF methods. Motivated by its strategic importance in 6G and its distinct advantages over legacy systems, this article systematically explores the challenges, enabling technologies, and architectural frameworks necessary to realize intelligent multimodal ISAC.
@@ I-B Contributions
    Unified cloud-edge-terminal Framework: We present a unified Cloud-Edge-Terminal (CET) framework that distributes intelligence hierarchically. Specifically, we design three dynamic operational modes: (i) Global Fusion Mode (GFM), utilizing vertical links for high-precision centralized fusion; (ii) Cooperative Relay Mode (CRM), employing edge nodes as semantic bridges for collaborative guidance; and (iii) Peer Interaction Mode (PIM), facilitating horizontal interactions for agile autonomy. This sequence reflects a continuum of progressive decentralization, effectively shifting the computational locus from the global cloud to the cooperative edge, and finally to the autonomous terminal.
@@
    Case Validation: We validate the proposed CET framework through a comprehensive case study on the DeepSense 6G dataset [1]. The experimental results quantify the distinct advantages of each mode: the GFM maximizes accuracy, the PIM minimizes latency (offering a $\sim 7\times$ speedup), and the CRM effectively bridges the gap, offering the best trade-off between performance and latency. This shows the necessity of a mode-adaptive framework to meet the requirements in different scenarios.
@@ II-A Heterogeneous Multimodal Fusion
Multimodal ISAC inherently involves the integration of data streams originating from diverse sensing modalities such as RF signals, vision sensors, and LiDAR. These modalities differ substantially in terms of data dimensionality, spatial-temporal resolution, coverage, and semantic abstraction. For instance, aligning 2D visual images with 1D radar waveforms is particularly challenging due to discrepancies in data structure, coordinate systems, and information content [7]. Moreover, inconsistencies in sensing frequency and perception latency further complicate synchronized fusion.
@@ II-B High Communication Overhead
Real-time sensing and decision-making in distributed ISAC systems require frequent and high-volume data exchanges between sensors, edge nodes, and centralized servers [10]. This results in considerable bandwidth consumption and elevated energy expenditure, particularly in wireless or resource-constrained environments. The issue becomes more pronounced in low-dynamic scenarios, such as static nighttime surveillance, where redundant or minimally informative data (e.g., successive identical video frames) continue to be transmitted, thereby wasting transmission resources without improving situational awareness.
@@ II-C Context-Aware Architectural Design
Designing a system architecture that can adapt to diverse operational environments and task requirements remains a fundamental challenge. Rigid deployment modes may offer simplicity in control and resource management, but often lack the flexibility needed to accommodate the heterogeneous demands of ISAC applications. For example, while centralized architectures benefit from powerful computing infrastructure, they introduce latency and dependency issues in time-critical or infrastructure-sparse settings. Conversely, purely distributed systems offer autonomy but often suffer from limited coordination and reduced global situational awareness [12].
@@ III-C Multi-Agent System
MAS [15] provides the structural foundation for enabling scalable, decentralized intelligence within the multimodal ISAC. In MAS, multiple autonomous agents, either software-based or embodied in hardware, interact with the environment, as well as each other, to collaboratively achieve task objectives. Each agent perceives local conditions, makes decisions, and executes actions, often through cooperative reasoning and intent-aware communication. As shown in Fig. 1(c), the core component of MAS includes four functional modules: a knowledge base for storing domain knowledge, a memory module for accumulating interaction history, a LAM, such as Deepseek and GPT, for reasoning and decision-making, and external tools for executing actions. During operation, an agent continuously interacts with the environment, stores relevant observations in memory, combines them with contextual knowledge from the knowledge base, and inputs the fused information into the LAM. The LAM then analyzes the situation and invokes appropriate tools (e.g., an AI or generic algorithm) to perform task-specific actions.
@@
Figure 2: The illustration of the proposed CET framework for multimodal ISAC.
IV Cloud-edge-terminal Collaborative Multimodal ISAC Architectures
To fully harness the potential of the aforementioned enabling technologies (e.g., LAM, SC, and MAS) in the multimodal ISAC system, we propose a unified CET framework, as illustrated in Fig. 2. This hierarchical design is motivated by the practical constraints of emerging 6G applications, such as autonomous driving and urban surveillance. In real-world deployments, individual terminal sensors often lack the computational power to run LAMs locally, while relying solely on centralized cloud processing introduces prohibitive latency and excessive backhaul traffic for high-volume multimodal data. By distributing intelligence across three layers, this architecture strikes a balance between perception accuracy and resource efficiency. The terminal layer consists of heterogeneous sensors (e.g., cameras and radars) equipped with lightweight SC encoders for initial semantic extraction. The edge layer hosts MAS controllers at roadside units or base stations to manage local coordination and low-latency task offloading. Finally, the cloud layer leverages the powerful LAM to maintain global environment models and perform deep reasoning.
To accommodate diverse sensing tasks and dynamic network conditions, as shown in Fig. 3, the proposed framework operates in three distinct modes: GFM, CRM, and PIM. This sequence represents a progressive decentralization of intelligence, shifting the computational locus from the global cloud (GFM) to the cooperative edge (CRM), and finally to the autonomous terminal (PIM). The system dynamically transitions between these modes to optimize the trade-off between sensing accuracy, latency, and coverage.
Figure 3: The illustration of the three modes in the CET framework.(a) GFM, (b) CRM, and (c) PIM.
IV-A Global Fusion Mode
@@
As illustrated in Fig. 3(a), the GFM corresponds to the vertical data flow within our collaborative architecture, adopting a centralized fusion strategy to integrate semantic information from heterogeneous sensors. Specifically, multimodal sensors at the terminal layer are equipped with lightweight edge agents. These agents serve as semantic encoders, extracting and compressing task-relevant semantics from local multimodal data. Then, the encoded compact semantic representations are transmitted vertically to the cloud layer over wireless channels, passing through the edge infrastructure. At the cloud layer, a powerful LAM serves as the central agent to align and fuse the received multimodal semantics to generate unified global representations. The fused features are then utilized for downstream applications such as global decision-making, collaborative sensing, and large-scale environmental modeling.
A typical application scenario for GFM is intelligent transportation, where cameras, LiDAR sensors, and RF devices are distributed across various intersections [5]. In this context, each terminal sensor, via its corresponding edge agent, extracts semantic information such as vehicle positions, motion trajectories, and Doppler shifts. These compressed semantic features are aggregated at a central LAM located at a cloud-based traffic control center. Here, global fusion is performed to achieve a comprehensive understanding of real-time traffic conditions across the entire city grid, thereby enabling intelligent traffic signal coordination and city-wide congestion mitigation.
@@
In the GFM, the deployment of a powerful LAM at the central cloud enables deep reasoning over globally aggregated data, significantly improving decision accuracy. Meanwhile, SC offloads bandwidth-intensive raw data transmission from the terminal to the cloud, ensuring efficiency even in bandwidth-limited environments. However, this mode relies heavily on the connectivity to the Cloud Layer and the computational availability of the central agent. Consequently, when communication links between terminal sensors and the cloud are disrupted, or when latency constraints are stringent, the system may suffer from reduced responsiveness. This dependency on centralized infrastructure necessitates a shift towards edge-assisted coordination, leading to the Cooperative Relay Mode.
IV-B Cooperative Relay Mode
Fig. 3(b) illustrates the CRM, which serves as the intelligent bridging mechanism within our architecture, shifting the operational focus from the cloud to the edge layer. In this mode, the edge layer (e.g., roadside units (RSUs)) acts as a semantic relay and coordinator between heterogeneous sensors at the terminal layer. Unlike the vertical upload in GFM, CRM utilizes the edge node as an intermediary to facilitate local cross-modal coordination. Terminal sensors equipped with moderate edge agents extract essential feature semantics and transmit them to a nearby relay node. The relay agent then interprets, repackages, and forwards these semantics, often translating them into control directives, to other terminal agents operating on different modalities. This enables a “sense-and-guide” workflow where one modality cues another without burdening the central cloud.
A prime example of CRM is the drone-assisted surveillance scenario. Here, a camera-equipped drone (Terminal A) employs its onboard agent to extract positional semantics from aerial observations. This semantic information is transmitted to an RSU-based relay agent. Referencing a lightweight knowledge base, the relay agent formulates a precise semantic directive (e.g., “focus radar on coordinates (x, y)”) and forwards it to a ground-based radar sensor (Terminal B). Upon receiving the directive, the radar’s agent dynamically adjusts its ISAC beam toward the specified coordinates. This cooperative mechanism avoids exhaustive scanning and significantly improves sensing efficiency by leveraging the wide field-of-view of vision to guide the precise ranging capabilities of RF.
Compared to the GFM, CRM reduces dependency on the central cloud and supports extended spatial coverage through multi-hop relaying via edge nodes. However, it still relies on an infrastructure-based intermediary (the relay) to perform semantic translation and coordination. In scenarios demanding ultra-low latency or where infrastructure is unavailable, even the delay introduced by edge relaying may be unacceptable. This drives the need for a direct, infrastructure-free approach, as realized in the Peer Interaction Mode.
@@ IV-B Cooperative Relay Mode
Aspect GFM (Cloud-based) CRM (Edge-based) PIM (Terminal-based)
Sensing Coverage Global / Macro-scale Regional / Zone-based Local Neighborhood / One-hop
Strengths High-accuracy inference; Strong multimodal fusion Balanced workload; Moderate communication efficiency High autonomy; Strong scalability; Robust resilience
Weaknesses Strong central dependency; Low fault tolerance Protocol dependency; Compromised global/local reasoning Cognitive inconsistency; High terminal computational burden
Communication Load High Moderate Low
Use Cases Intelligent transportation hubs Drone surveillance Industrial inspection sites
IV-C Peer Interaction Mode
Fig. 3(c) depicts the PIM, which represents the final stage of decentralization: the horizontal cooperative path solely within the terminal layer. Designed to maximize autonomy, this mode enables direct, peer-to-peer (P2P) semantic interaction among distributed sensors, bypassing both the central cloud and edge relays. Each terminal agent, driven by lightweight language models, independently processes local multimodal inputs and transmits only essential high-level semantic outcomes, such as alerts, decisions, or event summaries, directly to neighboring terminal agents. Subsequently, receiving agents perform local analysis and reasoning by leveraging this shared semantic context, internal memory, and localized knowledge bases.
A representative application scenario for PIM is industrial inspection, where cameras, vibration sensors, and acoustic sensors are deployed along a production line [11]. When a vibration sensor agent detects an anomaly and a camera agent simultaneously identifies a visual defect, they can directly exchange semantic messages via the horizontal link to collaboratively diagnose and localize faults in real time. This direct terminal-to-terminal collaboration eliminates the round-trip latency to the cloud or edge, ensuring rapid response times critical for safety-sensitive tasks.
By removing the reliance on upper-layer infrastructure (cloud or edge), PIM achieves the highest level of resilience and is particularly effective in bandwidth-constrained or infrastructure-deficient environments. Nonetheless, the absence of a centralizing brain (GFM) or a coordinating relay (CRM) introduces challenges regarding cognitive consistency. Data heterogeneity and differences in prior knowledge across distributed terminal agents may lead to divergent interpretations. For example, without a higher-level node to align discrepancies, a camera-based agent and an infrared-based agent may generate conflicting local decisions under varying environmental conditions. These limitations highlight that while PIM offers superior speed and autonomy, dynamic switching back to CRM or GFM is necessary when higher-order reasoning or global consensus is required.
To clearly illustrate the differences among the three modes, Table I provides a comparative summary of their key characteristics.